OpenStudy (anonymous):

Solve the triangle. A = 19°, C = 102°, c = 6 Choices are: B = 31°, a ≈ 18, b ≈ 15.8 B = 59°, a ≈ 2, b ≈ 5.3 B = 59°, a ≈ 18, b ≈ 15.8 B = 59°, a ≈ 18, b ≈ 5.3

OpenStudy (anonymous):

Use law of sines

OpenStudy (anonymous):

sin(102 degrees)/6 = sin(19 degrees)/ a

OpenStudy (anonymous):

well I know B=59 but I think I doing something wrong

OpenStudy (anonymous):

sin(102 degrees)/6 = sin(19 degrees)/ a 6/sin(102 degrees) = a/ sin(19 degrees) a = (6 * sin(19 degrees))/ sin(102 degrees) = 1.997